# Improving Rule-based Reasoning in LLMs via Neurosymbolic Representations

### v1

@@ Abstract
Large language models (LLMs) continue to face challenges in reliably solving reasoning tasks, particularly tasks that involve precise rule following, as often found in mathematical reasoning tasks. This paper introduces a novel neurosymbolic method that improves LLM reasoning by encoding hidden states into neurosymbolic vectors, allowing for problem-solving within a neurosymbolic vector space. The results are decoded and combined with the original hidden state, boosting the model’s performance on numerical reasoning tasks. By offloading computation through neurosymbolic representations, this method improves efficiency, reliability, and interpretability. Our experimental results demonstrate an average of 82.86% lower cross entropy loss and 24.50 times more problems correctly solved on a suite of mathematical reasoning problems compared to chain-of-thought prompting and supervised fine-tuning (LoRA), while at the same time not hindering the performance of the LLM on other tasks.
@@ 1 Introduction
In this paper, we introduce a novel method that extends the capabilities of LLMs by encoding their hidden states into structured symbolic representations. Unlike previous work focusing on token-level program synthesis, our approach directly integrates symbolic algorithms within the neural model by running them in a symbolic space derived from the LLM’s internal representations. This innovation bridges the gap between neural and symbolic reasoning by extracting inputs from the LLM’s hidden state and operating directly on a structured, interpretable representation of the problem.
@@
  • A Neurosymbolic Method for LLMs: This work represents a first step toward integrating symbolic reasoning into LLMs by using predefined symbolic rules as a proof of concept. We explore the ability of symbolic algorithms to operate within a symbolic space constructed from the LLM’s latent representations.
@@ 2.2 Sparse Autoencoders
Sparse autoencoders (SAEs) are a class of unsupervised learning methods designed to parse high-dimensional data, such as the hidden states (also called activations) of LLMs, into sparse, monosemantic components (Olah et al., 2020; Le et al., 2021). These components, often referred to as ”concepts,” are linearly combined to reconstruct the original input data (Elhage et al., 2021). SAEs have been used to identify which latent features in an LLM are active during specific tasks, enabling researchers to explore the internal representations of the model. Furthermore, SAEs can be used to steer LLMs by selectively amplifying or suppressing certain concepts, providing a powerful tool for interpretability and control.
@@ 4.1 Dataset
The dataset used in this study consists of mathematical problems designed to evaluate the LLM’s reasoning capabilities across a range of tasks. Each problem involves two randomly generated integers, both constrained to a maximum of three digits. This constraint ensures that the input numbers can be represented as single tokens in the LLM’s vocabulary.
To maintain consistency in tokenization and output representation, operations that could produce results exceeding three digits apply a modulo operation to truncate the solution. For instance, the product of $932$ and $152$ is reduced to the last three digits by computing $(932\cdot 152)\mod 1000=816$. Each problem is formatted as a natural language query, such as: ”What is $932$ times $152$ mod $1000$?”
The selected problem types include:
@@
    Bitwise AND: Convert $x$ and $y$ into bitstrings, compute the AND, and convert back to base 10,
@@
    Bitwise XOR: Convert $x$ and $y$ into bitstrings, compute the XOR, and convert back to base 10,
@@
    Bitwise OR: Convert $x$ and $y$ into bitstrings, compute the OR, and convert back to base 10,
@@
These problem types were chosen to represent a diverse set of mathematical operations, ranging from basic arithmetic to more complex logical operations. By selecting such a wide variety of problems, the dataset ensures broad evaluation of the LLM’s rule-based numerical reasoning abilities.
To train the encoder and decoder networks, we do not use the addition and integer division problem types. This exclusion is intended to measure the potential degradation in system performance after being trained on a different set of problems, relative to other supervised fine-tuning strategies discussed later in this section.
@@ 4.3 Encoding Hidden States
The second stage, after prompting, involves converting the hidden states of the LLM into neurosymbolic vector representations. For this purpose, we train a linear encoder network designed to map the hidden states recorded during the forward pass into neurosymbolic vectors that represent the problem’s key components: the two input numbers and the operation type. For problems involving mod $1000$ to truncate the final three digits, the $1000$ is not represented as an input number, but instead is tied to a problem type (e.g., multiplication problem types will always apply modulo $1000$ to the final answer). The symbolic vectors are structured using the framework described in Section 3.1. The encoder is trained using a root mean squared error (RMSE) loss, with the objective of minimizing the difference between the predicted and true symbolic vectors.
@@ 4.4 Decoding Neurosymbolic States
After training, both the encoder and decoder networks are included in the LLM (as shown in Figure 1) to assist in solving mathematical reasoning problems. The process begins by encoding the hidden state of the designated LLM layer into a neurosymbolic vector. This vector is then queried to determine the problem type, which dictates the selection of an appropriate rule-based Python function. If the queried problem type is not sufficiently similar to any the problem types encountered during training, the decoder is bypassed, and the LLM proceeds with its standard forward pass. Otherwise, the predefined rule-based function is applied to the extracted input values from the neurosymbolic vector, generating a new neurosymbolic representation containing the computed solution. This solution vector is then decoded back into an LLM-compatible hidden state via the decoder network, allowing the model to incorporate the computed result into its forward pass.
@@ 4.5 Comparisons to Other Methods
We compared the performance of our method to two other popular strategies for improving the mathematical reasoning capabilities of LLMs: zero-shot CoT reasoning and supervised fine-tuning via LoRA modules. These methods were selected as baselines because they represent two distinct paradigms: implicit reasoning through prompting and explicit task-specific fine-tuning.
Chain of Thought reasoning (Wei et al., 2022; Kojima et al., 2022; Wang et al., 2022) involves prompting the model to generate intermediate reasoning steps explicitly, rather than directly providing a final answer. This approach encourages step-by-step reasoning, which is particularly beneficial for solving complex mathematical problems that require multi-step calculations or logical deductions (Zhou et al., 2022). CoT has been shown to improve interpretability and correctness in reasoning tasks by enabling the model to break down problems into smaller, manageable components (Nye et al., 2021; Wei et al., 2022). CoT prompting can be implemented by including examples of detailed reasoning in the training dataset or through few-shot prompting during inference (Kojima et al., 2022). This strategy leverages the model’s inherent capabilities without requiring architectural modifications, making it efficient for a wide range of reasoning tasks.
@@ 5.1 Encoder and Decoder Performance
Table 1: Performance of Symbolic, Standard, CoT, and LoRA LLMs on Various Problem Types. Note that Addition and Integer Division problem types are testing problems
@@
Mod NS LLM 98.0 0.187
Standard LLM 57.0 2.878
CoT LLM 74.0 4.139
LoRA LLM 18.3 3.608
Mult. NS LLM 97.5 0.233
Standard LLM 0.0 9.571
CoT LLM 20.0 12.044
LoRA LLM 2.3 4.941
GCD NS LLM 94.4 0.302
Standard LLM 62.8 1.355
CoT LLM 94.0 0.78
LoRA LLM 89.8 0.487
LCM NS LLM 88.2 1.09
Standard LLM 2.0 7.080
CoT LLM 9.0 14.468
LoRA LLM 3.8 5.002
Square NS LLM 65.3 2.896
Mod Standard LLM 3.0 5.606
CoT LLM 22.0 12.21
LoRA LLM 3.0 5.187
Bitwise NS LLM 92.3 0.802
And Standard LLM 2.0 7.737
CoT LLM 3.0 13.08
LoRA LLM 20.3 3.252
Bitwise NS LLM 97.3 0.24
Xor Standard LLM 7.0 9.213
CoT LLM 1.0 16.59
LoRA LLM 2.5 5.865
Bitwise NS LLM 86.9 0.5
Or Standard LLM 7.0 9.658
CoT LLM 1.0 14.842
LoRA LLM 2.5 5.865
Addition NS LLM 99.7 0.038
Standard LLM 100.0 5.52e-05
CoT LLM 77.0 2.84
LoRA LLM 1.0 10.362
Integer NS LLM 98.1 0.081
Division Standard LLM 99.0 0.020
CoT LLM 94.0 0.955
LoRA LLM 51.0 2.465
5.2 Overall System Performance
The performance of the proposed Neurosymbolic LLM (NS LLM) is compared against three baseline models: a Standard LLM, a CoT reasoning LLM, and a LoRA fine-tuned LLM. Two metrics are used to evaluate model performance across various mathematical problem types:
  • Score (↑): The percentage of problems for which the model provides the correct answer with the highest probability.
  • Loss (↓): The categorical cross-entropy loss per problem, representing the negative log probability of the correct answer tokens.
Table 1 presents the detailed results. Key observations are summarized below.
Neurosymbolic LLM Performance
The Neurosymbolic LLM outperforms all baseline models across all trained problem types (i.e., all problem types except addition and integer division), achieving significantly higher scores and lower losses. For most problems, the loss is significantly reduced relative to the Standard LLM, and the accuracy approaches 100%.
However, on more complex tasks, such as LCM and square modulo, performance is slightly lower. This may be due to the complexity of the underlying forward-pass algorithm required for these problems, which makes applying interventions via a single decoder network more challenging. A potential improvement could involve using multiple decoder networks to insert neurosymbolic information at different stages of the forward pass, enabling more precise alignment with the LLM’s internal computations.
Another reason for the reduction in scores is the encoding error rate, which is the percentage of misclassified input digits. As shown in Figure 3, at layer 17, the errors for the hundreds, tens, and ones digit places are 1.5%, 2.0%, and 1.0%, respectively. Errors in generating the correct neurosymbolic representation of the input problems will result in an incorrect solution neurosymbolic vector, which increases the likelihood that the LLM outputs an incorrect response.
@@ 5.2 Overall System Performance
Baseline Model Comparisons
CoT LLM: The CoT LLM improves over the Standard LLM in tasks like GCD (94.0% score, 0.78 loss) and modulo (74.0% score, 4.139 loss). However, CoT performs worse on tasks like bitwise operations, with scores below 3.0%. This is likely due to the increased opportunity for errors in multi-step reasoning, such as incorrect bitstring conversion during intermediate steps. Furthermore, CoT strategies consistently exhibit higher loss values than other methods, reflecting the narrow token path required to generate correct outputs from reasoning steps.
LoRA LLM: While LoRA fine-tuning improves performance on some tasks, it underperforms on more complex operations and exhibits poor generalization to tasks it was not trained on, such as addition and integer division. This contrasts with the Neurosymbolic LLM, which adapts by avoiding interventions for unseen problem types, preserving its generality.

### v2

@@ Abstract
Large language models (LLMs) continue to face challenges in reliably solving reasoning tasks, particularly tasks that involve precise rule following, as often found in mathematical reasoning tasks. This paper introduces a novel neurosymbolic method that improves LLM reasoning by encoding hidden states into neurosymbolic vectors, enabling problem-solving within a neurosymbolic vector space. The results are decoded and merged with the original hidden state, significantly boosting the model’s performance on numerical reasoning tasks. By offloading computation through neurosymbolic representations, this method enhances efficiency, reliability, and interpretability. Our experimental results demonstrate an average of 88.6% lower cross-entropy loss and 15.4 times more problems correctly solved on a suite of mathematical reasoning tasks compared to chain-of-thought prompting and supervised fine-tuning (LoRA), while not hindering the LLM’s performance on other tasks. We make our code available at Neurosymbolic LLM [^1] https://github.com/vdhanraj/Neurosymbolic-LLM
@@ 1 Introduction
In this paper, we introduce a novel method that extends the capabilities of LLMs by encoding their hidden states into structured symbolic vector representations. Unlike previous work focusing on token-level program synthesis, our approach directly integrates symbolic algorithms within the neural model by running them in a symbolic space derived from the LLM’s internal representations. This innovation bridges the gap between neural and symbolic reasoning by extracting inputs from the LLM’s hidden state and operating directly on a structured, interpretable representation of the problem.
@@
  • A Neurosymbolic Method for LLMs: This work represents a first step toward integrating symbolic reasoning into LLMs . We explore the ability of symbolic algorithms to operate within a symbolic space constructed from the LLM’s latent representations.
@@ 2.2 Sparse Autoencoders
Sparse autoencoders (SAEs) are a class of unsupervised learning methods designed to parse high-dimensional data, such as the hidden states (also called activations) of LLMs, into sparse, monosemantic components (Olah et al., 2020; Le et al., 2021). These components, often referred to as “concepts,” are linearly combined to reconstruct the original input data (Elhage et al., 2021). SAEs have been used to identify which latent features in an LLM are active during specific tasks, enabling researchers to explore the internal representations of the model. Furthermore, SAEs can be used to steer LLMs by selectively amplifying or suppressing certain concepts, providing a powerful tool for interpretability and control.
@@ 4.1 Dataset
We release a formally specified, procedurally generated benchmark, the Symbolic‑Math Dataset [^2] https://github.com/vdhanraj/Symbolic-Math-Dataset, to foster reproducible evaluation of arithmetic reasoning in LLMs. The dataset is open-source (MIT license) and fully regenerable, enabling reproducibility and scaling to more complex queries of the same arithmetic form (i.e., operations over arbitrarily many digits).
Construction.
In this study, each example is built by (i) sampling two independent three‑digit integers ($x,y\in\{0,\dots,999\}$) and (ii) sampling a problem type $t$ from a fixed set of $p=10$ symbolic operations (listed below). To ensure every operand and result remains a single sub‑word token in LLaMA‑3, we mod‑reduce any outcome that exceeds three decimal digits: e.g. $(932\!\times\!152)\bmod 1000=816$. The instance is rendered as a natural‑language question such as
‘‘What is 932 times 152 mod 1000?’’
and paired with the numeric answer encoded as a single token. The problem types used in this study are:
@@
    Bitwise AND: $\mathrm{int}(\mathrm{bin}(x)\,\&\,\mathrm{bin}(y))$,
@@
    Bitwise XOR: $\mathrm{int}(\mathrm{bin}(x)\,\oplus\,\mathrm{bin}(y))$,
@@
    Bitwise OR: $\mathrm{int}(\mathrm{bin}(x)\,\lor\,\mathrm{bin}(y))$,
@@
Separate training, validation, and test splits are procedurally generated. The training and validation sets exclude addition and integer division, which are included only in the test set to evaluate out‑of‑distribution generalization.
Prompting format.
In our study, each test query is presented in a few-shot format with two in-context exemplars of the same problem type, preceding the target question. This consistent demonstration style encourages the model to learn the syntactic and arithmetic patterns of the task from examples alone, promoting the model to provide responses in a consistent and easy to evaluate format.
@@ 4.3 Encoding Hidden States
The second stage, after prompting, involves converting the hidden states of the LLM into neurosymbolic vector representations. For this purpose, we train a linear encoder network designed to map the hidden states recorded during the forward pass into neurosymbolic vectors that represent the problem’s key components: the two input numbers and the operation type (see Figure 1). For problems involving mod $1000$ to truncate the final three digits, the $1000$ is not represented as an input number, but instead is tied to a problem type (e.g., multiplication problem types will always apply modulo $1000$ to the final answer). The symbolic vectors are structured using the framework described in Section 3.1. The encoder is trained using a root mean squared error (RMSE) loss, with the objective of minimizing the difference between the predicted and true symbolic vectors.
@@ 4.4 Decoding Neurosymbolic States
After training, both the encoder and decoder networks are included in the LLM (as shown in Figure 1) to assist in solving mathematical reasoning problems. The inference process begins by encoding the hidden state of the designated LLM layer into a neurosymbolic vector. This vector is then queried to determine the problem type, which dictates the selection of an appropriate rule-based Python function. If the queried problem type is not sufficiently similar to any the problem types encountered during training, the decoder is bypassed, and the LLM proceeds with its standard forward pass. Otherwise, the predefined rule-based function is applied to the extracted input values from the neurosymbolic vector, generating a new neurosymbolic representation containing the computed solution. This solution vector is then decoded back into an LLM-compatible hidden state via the decoder network, allowing the model to incorporate the computed result into its forward pass.
@@
4.5 Computational complexity.
Although our method introduces additional neurosymbolic processing steps, its computational overhead is minimal. As detailed in Appendix C, the total runtime cost per forward pass through the neurosymbolic block is
$\Theta\!\bigl{(}d\,v+v\log v\bigr{)},$
which is independent of the sequence length $n$ and number of layers $L$, and is asymptotically dominated by the standard key–value cached transformer cost of $\mathcal{O}(L(n\,d+d^{2}))$. The space overhead is likewise modest, $\Theta(dv)$, which is negligible relative to the memory usage of the full LLM. This confirms that the neurosymbolic extension can be deployed efficiently without impacting scalability.
@@ 4.6 Comparisons to Other Methods
We compared the performance of our method to two other popular strategies for improving the mathematical reasoning capabilities of LLMs: zero-shot chain-of-thought (CoT) reasoning and supervised fine-tuning via LoRA modules. These methods were selected as baselines because they represent two distinct paradigms: implicit reasoning through prompting and explicit task-specific fine-tuning.
Chain-of-Thought reasoning (Wei et al., 2022; Kojima et al., 2022; Wang et al., 2022) involves prompting the model to generate intermediate reasoning steps explicitly, rather than directly providing a final answer. This approach encourages step-by-step reasoning, which is particularly beneficial for solving complex mathematical problems that require multi-step calculations or logical deductions (Zhou et al., 2022). CoT has been shown to improve interpretability and correctness in reasoning tasks by enabling the model to break down problems into smaller, manageable components (Nye et al., 2021; Wei et al., 2022). CoT prompting can be implemented by including examples of detailed reasoning in the training dataset or through few-shot prompting during inference (Kojima et al., 2022). This strategy leverages the model’s inherent capabilities without requiring architectural modifications, making it efficient for a wide range of reasoning tasks.
@@
5 Experiments
5.1 Evaluation Setup
We evaluate the proposed Neurosymbolic LLM (NS LLM) against three baselines: (i) a Standard LLM (frozen, with few-shot prompting), (ii) a LoRA-fine-tuned LLM trained on the same task corpus, and (iii) a CoT prompted LLM.
All models are evaluated on the Symbolic-Math Dataset described in Section 4.1. We use a procedurally generated split consisting of 20,000 training examples, 200 validation examples, and 2,000 test examples. The training set is used to fit model parameters, the validation set tracks accuracy during training, and the test set is used for final evaluation.
Each model is prompted using the same few-shot format: two in-context exemplars of the same problem type precede the target query, as detailed in Section 4.1. For all approaches, generation uses greedy decoding (temperature = 0).
We report two evaluation metrics:
  • Score (↑): The percentage of test examples for which the model assigns highest probability to the correct answer.
  • Loss (↓): The categorical cross-entropy loss on the target token, i.e., the negative log-likelihood of the correct answer.
All reported results are generated using single runs of each approach over the entire testing dataset.
5.2 Base LLM
The base LLM is evaluated using the same few-shot prompt format described in Section 5.1, with two in-context examples preceding each query. The model performs a single forward pass to generate its prediction for the final answer token.
We use the LLaMA 3.1 8B model for all experiments, following the inference procedure and key–value caching mechanism outlined in Grattafiori et al. (2024). The model weights are frozen during evaluation, and no additional fine-tuning is applied.
5.3 NS LLM
The NS LLM follows the methodology described in Section 4. At a designated layer, the hidden state corresponding to the most recent token is passed through a trainable encoder network to produce a neurosymbolic vector representation. This vector undergoes neurosymbolic processing, after which it is decoded back into the LLM hidden state space using a trainable decoder network.
To avoid erroneous interventions, the decoder’s output is only incorporated into the LLM’s hidden state when the model is confident that the encoded neurosymbolic vector correctly reflects the problem type. Specifically, we compute the dot product similarity between the extracted neurosymbolic vector and each problem type vector in the vocabulary, and apply the decoder output only if the highest similarity exceeds a threshold of 0.8 (justification for this threshold is provided in Appendix A). This gating mechanism prevents the neurosymbolic procedure from modifying the LLM’s internal state on unfamiliar or out-of-distribution tasks, thereby preserving performance on problems that lack an associated neurosymbolic algorithm. Further discussion of the performance of the NS LLM on out-of-distribution tasks is provided in Appendix B.
In this study, we intervene at layer 17, as it achieves the lowest encoder reconstruction loss (see Section 6.1). The dimensionality of the vector symbolic architecture (VSA) is fixed at 2048. The decoder output is combined with the original hidden state using a 50/50 linear mixture. The empirical justification for this mixing strategy is provided in Appendix D.
The encoder and decoder networks are initially trained for 1,000 epochs to ensure accurate neurosymbolic representations. Subsequently, the decoder is fine-tuned for one epoch using cross-entropy loss to align its outputs with the LLM’s internal expectations during inference.
5.4 LoRA
To ensure a fair comparison with the NS LLM, we implement a LoRA module with rank 2048—matching the dimensionality of the VSA used in the neurosymbolic method. This ensures both approaches have an equivalent number of trainable parameters. As with the NS LLM, the output of the LoRA module is mixed with the original hidden state at the intervention layer using a 50/50 weighted sum.
The LoRA module is trained for 1 epoch to match the fine-tuning stage of the NS LLM. Unlike the NS LLM, LoRA does not undergo a symbolic pretraining phase, as its encoder output is unconstrained. In contrast, the NS LLM explicitly enforces its encoder to produce structured VSA-style representations, enabling neuro symbolic querying and interpretation.
5.5 CoT
For the Chain-of-Thought (CoT) baseline, the LLM is not prompted with few-shot exemplars. Instead, its system prompt instructs it to "Always explain your reasoning step by step", encouraging it to perform structured reasoning autonomously. This setup ensures that the model generates its own intermediate steps rather than relying on algorithmic demonstrations embedded in the prompt.
@@ 6.1 Encoder and Decoder Performance
Table 1: Performance of Symbolic, Standard, CoT, and LoRA LLMs on Various Problem Types. Note that Addition and Integer Division problem types are not seen during training
@@
Mod NS LLM 98.7 0.093
Standard LLM 53.5 2.904
CoT LLM 69.7 4.424
LoRA LLM 51.5 3.838
Mult. NS LLM 95.6 0.314
Standard LLM 1.1 9.279
CoT LLM 25.3 11.755
LoRA LLM 4.5 6.279
GCD NS LLM 94.2 0.205
Standard LLM 62.6 1.31
CoT LLM 93.2 0.874
LoRA LLM 74.5 1.235
LCM NS LLM 87.3 1.051
Standard LLM 2.5 7.359
CoT LLM 10.8 14.778
LoRA LLM 2.0 5.941
Square NS LLM 58.9 2.818
Mod Standard LLM 7.0 5.054
CoT LLM 32.7 9.934
LoRA LLM 5.5 5.600
Bitwise NS LLM 91.2 0.755
And Standard LLM 2.7 7.152
CoT LLM 5.5 11.556
LoRA LLM 9.0 4.670
Bitwise NS LLM 99.4 0.094
Xor Standard LLM 6.7 10.606
CoT LLM 1.1 16.606
LoRA LLM 8.0 6.116
Bitwise NS LLM 97.6 0.093
Or Standard LLM 4.4 9.527
CoT LLM 7.8 12.423
LoRA LLM 10.5 5.046
Addition NS LLM 98.2 0.206
Standard LLM 100.0 0.000
CoT LLM 78.8 2.218
LoRA LLM 46.5 6.299
Integer NS LLM 97.4 0.066
Division Standard LLM 95.2 0.148
CoT LLM 94.3 0.709
LoRA LLM 72.0 1.797
6.2 Neurosymbolic LLM Performance
Across all trained problem types, the Neurosymbolic LLM achieves the best overall performance among all models, as shown in Table 1. It consistently attains higher accuracy and lower cross-entropy loss. For most problems, both the loss is significantly reduced and the accuracy is much higher than that of the Standard LLM.
However, on more complex tasks, such as LCM and square modulo, performance is slightly lower. This may be due to the complexity of the underlying forward-pass algorithm required for these problems (e.g., square modulo requires two-hop reasoning), which makes applying interventions via a single decoder network more challenging. A potential improvement could involve using multiple decoder networks to insert neurosymbolic information at different stages of the forward pass, enabling more precise alignment with the LLM’s internal computations.
Another reason for the reduction in scores is the encoding error rate, which is the percentage of misclassified input digits. As shown in Figure 3, at layer 17, the errors for each of the digits are all under 2%. Errors in generating the correct neurosymbolic representation of the input problems will result in an incorrect solution neurosymbolic vector, which increases the likelihood that the LLM outputs an incorrect response.
@@ 6.2 Neurosymbolic LLM Performance
6.3 Baseline Model Comparisons
CoT LLM: The CoT LLM improves over the Standard LLM in tasks like GCD (93.2% score, 0.874 loss) and modulo (69.7% score, 4.424 loss). However, CoT performs worse on tasks like bitwise XOR, where the score drops from 6.7% (Standard LLM) to 1.1%. This is likely due to the increased opportunity for errors in multi-step reasoning, such as incorrect bitstring conversion during intermediate steps (further discussed in Appendix F). Furthermore, CoT strategies consistently exhibit higher loss values than other methods, reflecting the narrow token path required to generate correct outputs from reasoning steps.
LoRA LLM: While LoRA fine-tuning improves performance on some tasks, it underperforms on more complex operations and exhibits poor generalization to tasks it was not trained on, such as addition and integer division. This contrasts with the Neurosymbolic LLM, which adapts by avoiding interventions for unseen problem types, thereby preserving its generality.